### v1

@@ 5.2 Results
5.2.1 Detection Experiment Results
The performance metrics for the toxicity detection task are summarized in Table 6. All configurations demonstrated strong overall accuracy, with both GPT-4o and Gemini achieving around 0.87 accuracy in the zero-shot setting. Notably, GPT-4o zero-shot exhibited balanced precision and recall for both classes, with scores around 0.86–0.87. In contrast, GPT-4o few-shot showed a trade-off: it achieved a higher Toxic recall (0.90) compared to zero-shot (0.86), but the improvement came at the expense of slightly lower overall accuracy (0.84) due to a decrease in Non-Toxic recall (0.79).
Gemini zero-shot performed well across both classes, achieving the highest Non-Toxic recall (0.92) and strong Toxic precision (0.90), indicating a more conservative but effective approach. However, Gemini few-shot exhibited a more aggressive stance towards detecting Toxic content, reaching the highest Toxic recall (0.97). This came at the expense of increased false positives, leading to a drop in overall accuracy (0.77) and a significantly lower Non-Toxic recall (0.58).
Compared to the previous study by Ramen, which used an SVM classifier on the same test set and achieved 0.75 precision and 0.35 recall for Toxic content, the LLM-based approaches significantly improved both precision and recall. This highlights the potential of LLMs in toxicity detection, offering a more balanced and effective classification of Toxic and Non-Toxic content. While few-shot prompting generally enhanced Toxic recall, it also led to an increase in false positives, suggesting the need for careful calibration to optimize performance across both classes.
Table 6: Evaluation of Prompting Methods in Toxicity Detection
@@ 5.2.1 Detection Experiment Results
GPT-4o zero-shot 0.87 0.87 0.87 0.87 0.86 0.86 0.86
GPT-4o few-shot 0.84 0.90 0.79 0.84 0.80 0.90 0.85
Gemini zero-shot 0.87 0.84 0.92 0.88 0.90 0.81 0.85
Gemini few-shot 0.77 0.95 0.58 0.72 0.68 0.97 0.80
Figure 7: Evaluation of Prompting Methods in Toxicity Detection
5.2.2 Mitigation Experiment Results
All toxic comments in our sample were successfully mitigated with a single rewrite, achieving a Mitigation Success Rate of 100.00%. The average number of attempts remained at 1.00, indicating that minimal rewriting was typically sufficient to transform a toxic comment into an acceptable form. Fig. 6 showcases examples of toxic comments alongside their rewritten versions. Upon examination of these examples, it becomes evident that LLMs can effectively remove toxicity while preserving the original meaning of the comments. Furthermore, for cases with relatively low similarity scores, we observed that these instances often contained little substantive content beyond their toxic elements. This suggests that when a comment is predominantly toxic, the rewritten version necessarily diverges more from the original.
5.3 Next Steps
@@
5.3.1 Domain-Specific Model Training
One key area for improvement is training LLMs on software engineering discourse. Prior research shows general-purpose models struggle with domain-specific terminology and context, leading to misclassifications [45]. Fine-tuning models using labeled SE datasets could enhance their ability to detect nuanced toxicity in developer communications.
5.3.2 Optimized Prompt Engineering
Our results suggest that prompt design significantly impacts the effectiveness of LLMs in both detection and mitigation. Previous studies have demonstrated that structured prompts can improve precision and recall in toxicity detection, as prompt clarity, specificity, and framing affect model performance [34]. Future work should explore advanced prompt engineering techniques, such as few-shot learning and chain-of-thought prompting, to optimize model outputs and minimize false positives or false negatives.
5.3.3 Explainable Toxicity Detection
While current approaches achieve high accuracy, they often lack interpretability, which is crucial in SE settings where developers require actionable insights. Research has highlighted the importance of integrating explanation mechanisms into LLM-based toxicity classifiers, such as providing rationale generation or highlighting specific toxic phrases [34]. Future work should investigate explainability frameworks that enhance model trust and usability while ensuring transparency in classification decisions.
@@ 6 Conclusion
In summary, this paper not only provides a comprehensive overview of current methodologies but also underscores existing limitations and the need for further refinement. Future research should focus on domain-specific model tuning, advanced prompt engineering, and enhanced interpretability to build more transparent toxicity mitigation systems. Through these efforts, we hope to pave the way for safer and more inclusive online environments.

### v2

@@ 6 Results
6.1 Toxicity detection performance
The performance metrics for the toxicity detection task across all code review comments are summarized in Table 13. Among zero-shot configurations, Gemini-1.5-flash zero-shot achieved the highest overall accuracy (0.85), closely followed by GPT-4.1 zero-shot (0.82). DeepSeek-chat zero-shot performed comparably accuracy(0.84) and showed the highest Toxic recall (0.54) and F1-score (0.56) among all zero-shot models, indicating stronger sensitivity to detecting toxic comments. All three zero-shot models All three models demonstrated strong Non-Toxic detection, with $precision\geq 0.83$ and $recall\geq 0.91$.
In contrast, few-shot prompting yields mixed results: GPT-4.1 few-shot reached 0.78 accuracy but suffered a drop in Toxic precision (0.40) with no improvement in Toxic recall (0.20) compared to zero-shot. Gemini few-shot improved Toxic recall to 0.55 (from 0.40) but at the expense of precision (0.33) and Non-Toxic recall (0.72), resulting in lower overall accuracy (0.69). DeepSeek few-shot similarly did not benefit from example-based prompting: while maintaining accuracy (0.83), Toxic recall declined to 0.42 and the corresponding F1-score dropped to 0.49.
Table 13: Evaluation of prompting methods in toxicity detection (full dataset)
@@ 6.1 Toxicity detection performance
GPT-4.1 zero-shot 0.82 0.83 0.97 0.90 0.67 0.20 0.31
GPT-4.1 few-shot 0.78 0.82 0.93 0.87 0.40 0.20 0.27
Gemini-1.5-flash zero-shot 0.85 0.87 0.96 0.91 0.73 0.40 0.52
Gemini-1.5-flash few-shot 0.69 0.87 0.72 0.79 0.33 0.55 0.42
DeepSeek-chat zero-shot 0.84 0.91 0.91 0.91 0.59 0.54 0.56
DeepSeek-chat few-shot 0.83 0.93 0.90 0.91 0.60 0.42 0.49
Table 14: McNemar’s test results for zero-shot vs. few-shot prompting across models.
Model $b$ $c$ $\chi^{2}$ $p$-value
GPT-4.1 447 331 16.999 <0.001<0.001
Gemini-1.5-flash 2321 634 961.961 <0.001<0.001
DeepSeek-chat (V3.2-Exp) 797 718 4.016 0.045
To assess the statistical significance of differences between zero-shot and few-shot prompting, we conducted McNemar’s test on the paired predictions of each model. Table 14 reports the off-diagonal counts ($b$: zero-shot correct, few-shot incorrect; $c$: zero-shot incorrect, few-shot correct), the test statistic, and the associated $p$-value. For both GPT-4.1 ($\chi^{2}=16.999$, $p<0.001$) and Gemini ($\chi^{2}=961.961$, $p<0.001$), the results indicate highly significant differences in overall error rates between zero-shot and few-shot settings. deepseek-chat also exhibits a statistically significant difference between prompting modes ($\chi^{2}=4.016$, $p=0.045$), but the effect size is much smaller, suggesting that DeepSeek’s detection performance is comparatively less sensitive to prompting strategy and that its strongest performance can be achieved without example-based conditioning.
Crucially, when comparing our results to supervised models, our best zero-shot LLM configuration (Gemini zero-shot) attains only 0.85 accuracy and a 0.50 Toxic F1 score. This is substantially outperformed by sarker2023automated’s supervised ToxiCR approach, which achieved 0.96 accuracy and an 0.89 Toxic F1 score on 19,651 code review comments. This suggests that domain-specific, supervised models still substantially outperform zero/few-shot LLM detectors on SE toxicity detection.
6.2 Toxicity mitigation performance
The mitigation experiment was applied to all comments initially flagged as toxic. Both Gemini-1.5-flash and deepseek-chat demonstrated strong mitigation capability, with overall success rates exceeding 99.5%. Gemini achieved a slightly higher Mitigation Success Rate (99.92%) and fewer failure cases (3 out of 3,547) compared to DeepSeek (99.52%, 7 failures). However, the two models exhibited different characteristics in terms of semantic preservation. While the average number of rewriting attempts was nearly identical for both models (1.02), DeepSeek produced rewritten comments with higher semantic similarity to the original text (mean similarity: 0.6251 vs. 0.5874 for Gemini), and also yielded a higher proportion of high-similarity rewrites (17.17% of outputs with similarity >0.9>0.9, compared to 9.13% for Gemini). This indicates that DeepSeek is more effective at maintaining the author’s original intent and technical nuance during toxicity reduction.
Fig. 6 presents representative examples of mitigated comments. In cases where similarity scores are lower for either model, we observe that the original comments often contained little meaningful content beyond their toxic phrasing. In such instances, preserving semantic equivalence is inherently more difficult because removing toxicity necessarily eliminates the primary information conveyed. Overall, these results suggest that while Gemini provides a more conservative and uniformly successful mitigation strategy, DeepSeek better preserves the communicative clarity and technical specificity that are especially important in code review contexts.
6.3 Study limitations
Internally, our analysis is constrained by the specific experimental characteristics of the ablation study. First, the limited sample size used in the LLM experiments restricts the statistical power of the analysis and may not fully capture the complete diversity of toxic expressions found in real-world SE communications. An analysis of cases where every LLM failed to detect toxicity reflects the inherent limitations of binary, generalized detectors when applied to developer communications: these include mild or implicit toxicity (e.g., “yeah that sucked, fixed. Done.”), judgmental critiques embedded in technical context (e.g., “function keyword is a bashism and IMHO looks ugly”), and the use of colloquial or informal language with mild profanity (e.g., ”confused the hell out of me”). Furthermore, the absence of extensive hyperparameter tuning and reliance on a single epoch of training for the fine-tuned supervised models mean that the performance differences reported may not represent the absolute optimal capacity of those supervised baselines. Similarly, the absence of k-fold cross-validation hinders our ability to fully assess the stability and robustness of the LLM and baseline model predictions across different partitions of the data. Readers should interpret the reported accuracy and F1 scores within the context of these resource and experimental design limitations.
Externally, our review is primarily constrained by a pervasive issue in SE toxicity research: the overwhelming availability of English-language datasets. Given that software development is an inherently global and multilingual process, relying exclusively on English data limits the broader applicability of these findings to development teams communicating in languages such as Chinese, Spanish, or Arabic. This linguistic bias means the identified model failures and mitigation successes may not translate effectively to non-English SE contexts, underscoring a necessary direction for future corpus expansion.
7 Next steps
@@
  • Continuous toxicity spectrum: We envision replacing the current binary classification with a continuous toxicity spectrum, assigning each comment a real-valued score between 0 and 1. This fine-grained approach would allow us to detect and quantify low-intensity toxic content that falls below rigid binary thresholds, addressing the issue of mild or implicit toxicity missed by current models. However, moving to a continuum demands more subjective annotation guidelines and requires careful calibration and validation to ensure consistency and reliability across evaluators.
  • Domain-specific model training: One key area for improvement is training LLMs on software engineering discourse. Prior research shows general-purpose models struggle with domain-specific terminology and context, leading to misclassifications [sarker2022built]. Fine-tuning models using labeled SE datasets could enhance their ability to detect nuanced toxicity in developer communications. Beyond supervised fine-tuning, future work should explore complementary strategies such as semi-supervised learning [zhu2005semi], leveraging unlabeled SE comments for pretraining to capture broader domain language patterns. Additionally, data augmentation techniques, like back-translation [edunov2018understanding] of rare toxic examples, could address challenges posed by low-resource toxicity classes, enriching training data and improving model robustness.
  • Optimized prompt engineering: Our results suggest that prompt design significantly impacts the effectiveness of LLMs in both detection and mitigation. Previous studies have demonstrated that structured prompts can improve precision and recall in toxicity detection, as prompt clarity, specificity, and framing affect model performance [mishra2024exploring]. Future work should explore advanced prompt engineering techniques, such as few-shot learning and chain-of-thought prompting, to optimize model outputs and minimize false positives or false negatives.
  • Explainable toxicity detection: While current approaches achieve high accuracy, they often lack interpretability, which is crucial in SE settings where developers require actionable insights. Research has highlighted the importance of integrating explanation mechanisms into LLM-based toxicity classifiers, such as providing rationale generation or highlighting specific toxic phrases [mishra2024exploring]. Future work should investigate explainability frameworks that enhance model trust and usability while ensuring transparency in classification decisions.
  • Phased Research Roadmap: We propose a phased research roadmap to guide the iterative implementation and evaluation of these advancements:
    • Phase 1: Corpus Expansion and Annotation (Estimated: 2-4 months). Focus on collecting and rigorously annotating a larger, diverse software engineering toxicity corpus from various SE platforms, capturing domain-specific nuances and low-resource toxicity classes, including developing guidelines for the continuous toxicity spectrum.
    • Phase 2: Model Adaptation and Optimization (Estimated: 1-2 months). Build upon the expanded corpus by fine-tuning existing LLMs or developing new domain-specific architectures. This involves exploring training strategies like semi-supervised learning and advanced prompt engineering to optimize model performance.
    • Phase 3: Explainable Prototype Deployment and User Evaluation (Estimated: 1-2 months). Deploy an explainable prototype toxicity detection system in a controlled SE environment, integrating explainability frameworks for actionable insights. Refine model performance using Reinforcement Learning from Human Feedback (RLHF, christiano2017deep) based on user preferences and context. Conduct thorough user evaluations to gather feedback, assess usability, and refine the system for real-world applicability.
@@ 8 Conclusion
Beyond technical performance, the implications of improved toxicity detection and mitigation in SE extend significantly to organizational culture [bhat2021say] and developer well-being [sarker2022identification]. By fostering healthier, more inclusive open-source communities, these advancements can enhance collaboration, improve developer retention, and support a more equitable and welcoming environment for all contributors. Ultimately, reducing toxic interactions can lead to more productive and sustainable development ecosystems.
In summary, this paper not only provides a comprehensive overview of current methodologies for combating toxic language but also underscores existing limitations and the need for further refinement. Future research should focus on domain-specific model tuning, advanced prompt engineering, and enhanced interpretability to build more transparent toxicity mitigation systems. Through these efforts, we hope to pave the way for safer and more inclusive online environments.